### v1

# Controlling Impact Noise Propagation With Ceramic Metatiles

Abstract
We examine the application of metamaterials for impact noise insulation in tiled floors. A ceramic metatile is developed which makes use of phononic metamaterials and optimized joint configurations. First, we optimize the geometrical and material parameters of the proposed metatile, which is composed of small ceramic tiles connected by silicon joints, in order to reduce longitudinal and flexural wave propagation on tiled floors. A bandgap is achieved that effectively mitigates the transmission of impact noise through periodic structural pathways. In both experimental and numerical tests, it is demonstrated that the integration of silicon joints inside the ceramic metatile improves the acoustic insulation performance, as measured by the reduction of impact noise levels across a wide range of low frequencies. The findings highlight the potential of metamaterials in architectural acoustics, offering innovative solutions for sound control in tiled environments.
I Introduction
One of the most significant factors in building design is to provide a comfortable environment for its residents. Several elements influence a home’s degree of comfort, including temperature, humidity, noise, air quality and hygiene. Neglecting such living conditions will have a negative impact on household comfort, e.g., noise generated by footsteps and noise pollution from surrounding buildings, temperature rise due to heat absorbed by buildings during the daytime, inadequate sanitation, and poor indoor air quality [1, 2]. Among these living conditions, noise control is a critical aspect of architectural design and urban planning, with far-reaching implications for physical and mental health. Reducing noise in enclosed spaces is a long-standing objective in architectural acoustics, especially in contexts like tiled rooms, where sound reflections and transmissions are enhanced by hard, reflective surfaces [3]. In general, there are numerous ways of controlling noise annoyance both within and outside buildings. Typical noise mitigation methods, such as the use of porous absorbers or heavy insulation, are often encountered with aesthetic, functional, or structural constraints. Recent advances in metamaterials [4, 5, 6, 7, 8, 9], including phononic crystals and acoustic metasurfaces, have opened new avenues for sound control in buildings, enabling the manipulation of wave propagation using compact and tunable designs [3, 10]. Thus, the solution to improve the comfort of noise insulation in building floors should be based on phononic metamaterials in one part and institutive materials in the other part.
Phononic metamaterials, which consist of periodic arrangements of scatterers and resonators with contrasting acoustic properties, are particularly promising due to their ability to create low frequency band gaps that prevent propagation of specific wave modes [11, 12, 13]. These bandgaps arise from Bragg scattering and local resonances, which can be engineered to effectively control longitudinal, transverse, and flexural polarized wave propagation. Such characteristics make phononic metamaterials perfect candidates for application in structural elements (e.g., ceramic tiles) to cope with the challenges of noise control in architectures [14, 15, 16, 17, 18]. Numerous studies have explored the potential of phononic crystals for vibration and noise control. For example, Maldovan demonstrated the tunability of acoustic bandgaps in phononic crystals, highlighting their role in controlling low-frequency noise [14]. Similarly, Hussein et al. reviewed the mechanisms of wave attenuation in phononic structures, highlighting their applicability in multi-scale designs for noise control [15]. In the context of architectural applications, research has focused on integrating phononic crystals into building components. Zhao et al. showed how periodic structures embedded in panels could attenuate noise transmission effectively [16], while Ghaffarivardavagh et al. utilized acoustic metamaterials to achieve low-frequency sound attenuation in compact spaces [19]. Another critical aspect of sound control is absorption. Traditional absorptive materials rely on viscous and thermal dissipation mechanisms, but recent studies have shown that metamaterials can achieve efficient sound absorption through localized resonances and impedance matching [20, 21, 22]. For instance, Mei et al. demonstrated that hybrid structures combining phononic crystals and acoustic absorbers could enhance overall performance [20]. Further advancements introduced broadband absorptive designs using subwavelength resonators, paving the way for efficient noise mitigation in compact configurations [22, 23].
Ceramic tile is a widely used construction material which is applied in various parts of buildings, including household roofs, walls, floors, and partitions [24]. In modern architectures, floor tiles are often designed for durability, resistance to abrasion, water impermeability, ease of cleaning, and visual appeal [25]. However, when it comes to acoustic properties, ceramic tiles are not ideal soundproofing materials. Their high density and hard surface make them less effective in reducing sound transmission compared to materials such as wood, laminate, or rubber. Targeting on noise insulation in tiles, Kumar et al. proposed a metatile prototype based on a circular maze-like acoustic metastructure, utilizing a space-coiling technique to enhance acoustic performance in the low-to-mid frequency range [26]. Choi et al. investigated locally resonant metamaterials applied to floors in under-construction residential buildings, conducting an in-depth study on vibro-acoustic control of floor impact sound [27]. Lopes et al. developed a methodology for maximizing broadband impact sound insulation by designing slabs with a non-uniform material thickness distribution, which was numerically optimized [28]. Finally, Sukontasukkul et al. explored the use of viscoelastic polymer sheets as an acoustic control treatment integrated into ceramic tiles to enhance their sound insulation properties [29]. Despite of these significant advancements, achieving effective noise insulation in ceramic tiles remains challenging, particularly in developing an optimal design specifically suited for tiled flooring applications. Previous works have established the role of tile size and material properties in tuning the bandgap frequencies for such applications [30, 16]. Additionally, silicon joints are utilized as coupling elements to optimize wave transmission and reflection, aligning with findings by Fang et al. [30] on the effectiveness of elastomeric joints in acoustic metamaterials.
Building on these principles, this research explores the use of phononic metamaterials and silicon joint design to reduce impact noise in tiled rooms. Inspired by the concept of metatiles [26], a term combining metamaterials and construction tiles, we propose a new design that integrates ceramic plates and silicon. This approach preserves the key properties of ceramic tiles, such as durability, abrasion resistance, water impermeability, ease of cleaning, and aesthetic appeal, while also improving their noise reduction capabilities. First, we analyze and optimize the geometrical parameters and material properties of metatiles, interconnected by silicon joints, to suppress longitudinal and flexural wave propagation. Next, we calculate the dynamic response using the finite element method. Finally, we present experimental validation and evaluate the acoustic performance of ceramic metatiles in impact noise insulation, demonstrating their potential for practical applications in architectural acoustics for tiled floors.
Figure 1: Schematic illustration of the proposed ceramic metatile. (a) Overview of the design strategy for impact noise insulation on ceramic metatiled floor. (b) Isometric view of the ceramic metatile with geometrical parameters. (c) Numerical model for transmission calculation, where the green dot represents the input displacement and the red dot represents the output displacement: (i) along the $[100]$ direction, (ii) along the $[110]$ direction, and (iii) along the $[010]$ direction.
II Theoretical Dispersion and Transmission
II.1 Design of the ceramic metatile
The proposed ceramic metatile is a periodic structure composed of ceramic sub-tiles and silicon joints. Fig. 1(a) shows an isometric view of the main principle behind using ceramic metatiles to mitigate sound propagation from impact noise or footstep noise. As depicted, the metatile floor consists of ceramic metatiles adhered to reinforced concrete using an adhesive. The periodic arrangement of ceramic sub-tiles and silicon joints resembles a monoatomic chain of resonators, with the silicon joints acting as springs and the ceramic sub-tiles serving as masses. This design enables sound wave control through the periodic mass-spring system, using Bragg scattering and local resonance between neighboring sub-tiles. The geometric parameters of the metatile are illustrated in Fig. 1(b), where $a$ represents the periodicity constant of a standard ceramic tile, $c$ is the width of the silicon joints, $b$ is the width of the sub-tiles, and $e$ is the thickness of the ceramic tiles taken as $e=1\,\text{cm}$ in this study. The standard ceramic tiles, provided by IRIS, were characterized through static and dynamic tests under small strains, yielding following estimates for their elastic properties: Young’s modulus $E_{c}=55\pm 5\,\text{GPa}$, Poisson’s ratio $\nu_{c}=0.3$, and density $\rho_{c}=3000\pm 300\,\text{kg/m}^{3}$. For the silicon material, we use a commercially available silicon with the following mechanical properties: Young’s modulus $E_{s}=1\text{MPa}$, Poisson’s ratio $\nu_{s}=0.4$, and density $\rho_{s}=1000\text{kg/m}^{3}$.
II.2 Numerical models
To characterize the noise impact reduction of our structure, we conduct a series of numerical simulations using the finite element method in the commercial software COMSOL Multiphysics. Our goal is to investigate the effect of adding silicon joints to the standard tiles on impact and footstep noise insulation. In doing so, we account for the linear elastic properties of the material by implementing a numerical model based on the weak formulation of the time-harmonic Navier equation, as follows:
$\frac{E}{2(1+\nu)}\biggl{(}\frac{1}{(1-2\nu)}\nabla(\nabla.\textbf{u})+\nabla^{2}\textbf{u}\biggl{)}=-\omega^{2}\rho\textbf{u},$ (1)
where $\textbf{u}=(u,v,w)$ is the displacement vector, $\omega$ is the angular wave frequency and $\nabla$ is the gradient, $E$ is the Young’s modulus, $\nu$ is the Poisson’s ratio.
First, we compute their phononic dispersion relations, assuming that the ceramic metatiles are repeated periodically along the $x$-direction and finite in the $y$-axis. We apply the Floquet-Bloch boundary conditions along the $x$-direction as:
$\textbf{u}(x+d)=\textbf{u}(x)e^{ikd},$ (2)
where ${\bf k}=(k,0,0)$ is the reduced Bloch wavevector and $\textbf{u}(x)$ is the displacement vector. The cells are assumed to be periodically repeated along the $x$-direction with periodicity $d=a+5c$.
The phononic dispersion relations are calculated in the first irreducible Brillouin zone $(\Gamma X)$ in the reciprocal space, where $\Gamma=(0,0,0)$ and $X=(\pi/d,0,0)$. For the meshing, we take into account a maximum element size of $\lambda_{\text{min}}/10$ and tetrahedral elements.
Following our examination of the phononic dispersion in ceramic metatiles (infinite), we proceed to investigate their transmission characteristics through simulations involving a single metatile (finite). We compare the results for three different directions for both the standard tile and the metatile, as depicted in Fig. 1(c). The transmission of the elastic waves is calculated using the following equation:
${\rm Transmision}\;{(\rm dB)}=20\log_{10}\left(\frac{\sqrt{|u|^{2}+|v|^{2}+|w|^{2}}_{\rm output}}{\sqrt{|u|^{2}+|v|^{2}+|w|^{2}}_{\rm input}}\right),$ (3)
where $\sqrt{|u|^{2}+|v|^{2}+|w|^{2}}_{\rm input}$ and $\sqrt{|u|^{2}+|v|^{2}+|w|^{2}}_{\rm output}$ are the modulus of the displacement vectors of the input and output elastic waves, respectively.
II.3 Numerical results
The main focus of this study is the investigation of bandgaps, which indicates frequency ranges where the ceramic metatile effectively attenuates footstep and impact noise vibrations. To ensure accurate and reliable results, it is essential to establish a methodology for determining the geometrical dimensions and silicon parameters of the proposed ceramic metatile. We start by analyzing the bandgaps extracted from the phononic dispersion in the first irreducible Brillouin zone, specifically along the $\Gamma X$ direction, for a unit cell consisting of two subtiles of width $b$ and two silicon joints of width $c$. The bandgaps variations are examined as functions of the geometrical parameters of the ceramic metatile: the width of the subtiles $b$ and the width of the silicon joints $c$. Additionally, we investigate the influence of the Young’s modulus of silicon on these bandgaps. The eigenvalue calculations are presented in Fig. 2 for an infinite structure along the $[100]$ direction and $[010]$ direction.
Firstly, we fixed the width of the silicon joints at $c=0.5$ cm and the Young’s modulus of silicon at $1$ MPa. Then, we varied the parameter $b$ from $2$ cm to $10$ cm to analyze its effect on the bandgaps. Fig. 2(a) presents the calculated bandgaps in function of $b$, which are colored in gray, along with their corresponding frequencies. It is observed that for $b=2$ cm, the bandgaps start from $506$ Hz and extend up to approximately $1865$ Hz. As the width of the subtiles increases, the bandgaps shift toward lower frequencies, reaching a minimum frequency of $232$ Hz. This indicates that when the subtiles are larger, bandgaps appear in the lower frequency range. However, as their size continues to increase, the diffraction limit shifts further to lower frequencies, ultimately disrupting the bandgap around $1500$ Hz.
[…]
Secondly, for the variation of the silicon joint width ($c$), we fixed the width of the subtiles at $b=5$ cm and the Young’s modulus of silicon at $1$ MPa, as depicted in Fig. 2(b). Then, we varied $c$ from $0.1$ cm to $0.5$ cm to limit the size of silicon joints, ensuring that the ceramic metatile retains its resemblance to a standard one, with only small additional silicon joints. It is observed that for $c=0.1$ cm, the bandgap starts at $831$ Hz. As the width of the silicon joints increases, this bandgap shifts toward lower frequencies, reaching $328$ Hz for $c=0.5$ cm. This parametric study provides insights into how the silicon joints influence the frequency range of the bandgaps and how they can be controlled by introducing silicon joints into the ceramic metatile. However, the dimensions of the silicon joints also depend on the target frequency range for the intended application.
In the final parametric study, we fixed the dimensions of the ceramic metatile with a subtile width of $5$ cm and a silicon joint width of $0.5$ cm, as depicted in Fig. 2(c). We varied the Young’s modulus of silicon to better understand how silicon mechanical properties affect noise cancellation. By keeping the density and Poisson’s ratio of the silicon constant, we varied the Young’s modulus from $1$ MPa to $10$ MPa. It is observed that as the Young’s modulus decreases, the bandgaps shift to lower frequencies. For example, for $E=1$ MPa, the bandgaps start at $328$ Hz. As the silicon becomes stiffer, the bandgaps shift to higher frequencies. For $E=10$ MPa, the bandgaps start at $1037$ Hz.
Figure 3: Phononic dispersion of the optimized ceramic metatile in the first irreducible Brillouin zone ($\Gamma X$). The geometric and material parameters used are $E=1MPa$, $c=0.5cm$, and $b=5cm$.
This analysis of the material and geometric parameters provides us with a better understanding of the bandgaps and their appearance in the ceramic metatile. After this detailed analysis, we chose $c=0.5$ cm, $b=5$ cm, and a silicon Young’s modulus of $1$ MPa, which we consider for the rest of the paper. Using these values, we plot the dispersion curves in the first irreducible Brillouin zone along the $[100]$ direction. The results of this study are presented in Fig. 3. This figure shows the bandgaps, represented by the gray-colored region, and the propagating modes, which are plotted in blue. Additionally, we display the eigenmodes of the ceramic metatile around the bandgaps, shown as screenshots of the total modulus displacement distribution for a ceramic metatile composed of $3$ unit cells.
Figure 4: Numerical transmission curves for transverse waves. (a) Normal to the silicon joints along the $[100]$ direction. (b) Diagonal to the silicon joints along the $[110]$ direction. (c) Parallel to the silicon joints along the $[010]$ direction.
After calculating the phononic dispersion of the ceramic metatile, modeled as an infinite structure, we now proceed to characterize its dynamic behavior by analyzing the transmission of elastic waves through a single ceramic tile. This tile consists of $5$ silicon joints and $6$ ceramic subtiles, as illustrated in Fig. 1(c). The transmission spectrum is calculated along the three principal directions: $[100]$, $[110]$, and $[010]$, to assess the ceramic metatile’s response to wave propagation in different orientations.
Figure 5: Numerical transmission curves for longitudinal waves. (a) Normal to the silicon joints along the $[100]$ direction. (b) Diagonal to the silicon joints along the $[110]$ direction. (c) Parallel to the silicon joints along the $[010]$ direction.
We begin by applying a transverse wave and calculating its transmission, as illustrated in Fig. 4. Fig. 4(a) presents the computed transverse transmission along the $[100]$ direction, where the red curve corresponds to the ceramic metatile, the blue curve represents the standard ceramic tile and the black curve illustrates the transmission through one silicon joint. In this direction, the transmissions of the standard ceramic tile and the $1$ silicon joint remain high across the entire frequency range studied ($20-5000$ Hz). In contrast, the ceramic metatile exhibits a significant reduction in transmission within the $500-1900$ Hz range, whereas in the lower range ($20-500$ Hz), the transmission remains nearly total. Beyond $1900$ Hz, the bandgaps are affected by diffraction limits, yet they retain their general shape and continue to attenuate vibrations effectively.
In the $[110]$ direction (Fig. 4(b)), the bandgap extends from $500$ Hz to $1400$ Hz, after which the transmission increases until $2000$ Hz. Beyond this point, additional bandgaps appear, influenced by the diffraction limit. Meanwhile, the standard ceramic tile maintains nearly maximal transmission across the entire frequency spectrum. When analyzing transmission in the $[010]$ direction (Fig. 4(c)), where wave propagation is parallel to the silicon joints, we observe that the ceramic metatile’s transmission closely resembles that of the standard ceramic tile. This indicates that the presence of silicon joints in the propagation direction of transverse waves attenuates vibrations similarly to the inclined $45\textdegree$ joints. These findings suggest that our design effectively reduces transverse waves propagation in directions where the silicon joints are not aligned parallel to the wave propagation.
To further validate the dispersion curves in the first irreducible Brillouin zone for all existing modes, we now compute the transmission for a longitudinally polarized wave. Using the same model, we apply a longitudinal wave as the input displacement and measure the output displacement. The results of this study are illustrated in Fig. 5, where the transmission of the longitudinal wave is analyzed in three different directions. Fig. 5(a) shows the transmission along the $[100]$ direction. The standard ceramic tile exhibits excellent transmission of longitudinal vibrations, whereas the ceramic metatile displays significantly reduced transmission between $400$ Hz and $1950$ Hz. Beyond this range, bandgaps emerge, coupled with diffraction limits. Furthermore, we analyze the behavior of the inclined joints in relation to longitudinal wave propagation along the $[110]$ direction. Our design demonstrates a broad bandgap extending from $400$ Hz to $1950$ Hz, followed by the presence of bandgaps at higher frequencies, again influenced by diffraction effects, ad shown in Fig. 5(b). Finally, in the $[010]$ direction, the transmission behavior closely resembles that observed for transverse waves. In this case, both the standard ceramic tile and the ceramic metatile exhibit high transmission for both longitudinal and transverse waves, as depicted in Fig. 5(c).
We numerically demonstrate that our design, which incorporates silicon joints into ceramic metatiles, exhibits broadband bandgaps at low frequencies. Moreover, these bandgaps can be controlled by adjusting the geometrical and material parameters. The dispersion calculations presented in this section confirm that our chosen parameters create four distinct bandgaps in the low-frequency range of $20-2000$ Hz for an infinitely periodic structure. To further validate the presence of these bandgaps in a finite structure, we calculated the transmission of both transverse and longitudinal waves through a single ceramic metatile. Additionally, we numerically compared its response to that of a standard ceramic tile. In the next section, we present the experimental validation of these bandgaps.
III Experimental validation
III.1 Laboratory transmission measurements
Figure 6: Illustrations of lab transmission measurements and sample fabrication. (a) A signal generator is used to shape the incident field before getting amplified and sent to the piezo patches. The signal is then traveling as an elastic wave and thus collected (read) by an accelerometer. Same accelerometer is glued on the piezo patch to read the truly generated elastic oscillation. The data are collected via a signal analyser and stored on the computer (b) The standard ceramic tile used in our experiments. (c) The ceramic metatile fabrication procedure, which involves silicon joint structures.
To validate numerical results, a ceramic metatile was fabricated by assembling ceramic plates with silicon joints, as shown in Fig. 6(c). The fabricated metatile has geometric parameters of width of $32.5cm$ and $30cm$, with a height of $1cm$, which is composed of small ceramic plates of $5cm$ as width and separated by silicon joint with a width of $0.5cm$.
Fig. 6(a) illustrates the experimental setup used to measure the transmission spectrum of the ceramic metatile. A signal generator was employed to produce a sweep signal, which was fed into a power amplifier to enhance the signal amplitude. The amplified signal was then transmitted to a PZT patch, which generated the input displacement. A transducer measured this displacement, while another transducer at the output recorded the transmitted signal. The output transducer was connected to an analyzer, which interfaced with a laptop for data acquisition. This setup enabled the measurement and recording of both transmitted and incident displacements, allowing for the calculation of transmission curves.
Figure 7: Experimental transmission curves for transverse waves. (a) Normal to the silicon joints along the $[100]$ direction. (b) Diagonal to the silicon joints along the $[110]$ direction. (c) Parallel to the silicon joints along the $[010]$ direction.
III.2 Experimental results
For the experimental transmission measurements, we analyze the propagation of transverse waves along three directions: $[100]$, $[010]$, and $[110]$. Then, we compare the results obtained from a standard ceramic tile with those of our designed ceramic metatile. Fig. 7 presents the experimental results, where the blue curves correspond to the standard ceramic tile, and the red curves represent the ceramic meta-tile’s response in all directions. In the $[100]$ direction (Fig. 7(a)), the transmission remains high between $20$ Hz and $700$ Hz. Beyond this range, the transmission of the metatile decreases by approximately $50$ dB compared to the standard tile from $700$ Hz to $2500$ Hz. For the $[110]$ direction (Fig. 7(b)), a bandgap appears between $700$ Hz and $1400$ Hz. The transmission then increases until $1490$ Hz, after which the metatile exhibits a transmission reduction of around $50$ dB relative to the standard ceramic tile. finally, for the $[010]$ direction (Fig. 7(a)), both tiles exhibit high transmission of transverse waves.
IV Discussion
Figure 8: Impact noise application using a shaker along the [100] direction. (a) Standard ceramic tile subjected to an impact generated by a shaker. (b) Time response of the standard ceramic tile: the blue curve represents the measured impact noise source, while the red curve corresponds to the transmitted noise on the opposite side of the tile. (c) Ceramic metatile subjected to an impact generated by a shaker. (d) Time response of the ceramic metatile: the blue curve represents the measured impact noise source, while the red curve corresponds to the transmitted noise on the opposite side of the metatile. (e) and (f) FFT analysis of the source and transmitted vibrations for the standard ceramic tile and the ceramic metatile, respectively.
After numerically and experimentally demonstrating the existence of stop bandgaps in the ceramic metatile, this section focuses on studying impact noise reduction in both the standard ceramic tile and the metatile using a real impact noise source. The setup, illustrated in Fig. 8(a) and (c), involves a shaker connected to a hammer-like object to simulate the impact noise generated by footsteps on tiled floors. We excite the shaker by a theoretical signal, which is a Gaussian-like signal with an interval of $100$ ms and we employ the same laboratory setup as depicted in Fig. 6.
A transducer is placed to measure the impact noise source, while another transducer is placed on the opposite side of the tiles in order to measure their transmission noise amplitude. The results of this study are shown in Fig. 8 (b) and (d), which show the time-dependent acceleration amplitude for both the standard ceramic tile and the metatile. The impact noise source is represented by the blue curve, while the transmitted noise is shown in red. Standard ceramic tile exhibits high transmitted noise, characterized by a large amplitude over time. In contrast, the metatile shows a significantly lower amplitude, indicating its effectiveness in attenuating noise propagation.
To further analyze this behavior, the Fast Fourier Transform (FFT) of both responses is calculated and presented in Fig. 8(e) and (f). The standard ceramic tile allows noise transmission across the frequency spectrum from $20$ Hz to $2000$ Hz. However, the metatile demonstrates a significant reduction in noise transmission at low frequencies, and beyond $500$ Hz, no impact noise transmission is observed. These results highlight the strong noise control and attenuation capabilities of the metatile, making it a promising solution for tiled flooring applications.
V Conclusion
In conclusion, we have proposed a novel design of metatiles that combines ceramic tiles with silicon joints. This design highlights the essential properties of ceramic tiles in noise reduction as tiled floors. Broadband bandgaps at low frequency ranges are obtained for both transverse and longitudinal waves. For transverse waves, the ceramic metatile exhibits a significant reduction in transmission within the $500-1900$ Hz range along the $[100]$ direction and within the $500-1400$ Hz range along the $[110]$ direction, respectively. For longitudinal waves, the ceramic metatile displays a broad bandgap extending from 400 Hz to 1950 Hz along the $[100]$ and $[110]$ directions. Besides, the bandgaps shift towards lower frequencies as the width of subtiles and silicon joints increases, or as the Young’s modulus of silicon decreases, which indicates that they can be tailored by geometry parameters and material properties. Furthermore, we fabricate metatile samples and experimentally validate their strong noise control and attenuation capabilities, showcasing their potential for practical use in architectural acoustics for tiled floors.
Acknowledgements
This research was funded by IRIS, BTU Trust, the French ANR (contract "ANR-21-CE33-0015") and the European Union under Marie Skłodowska-Curie Actions Postdoctoral Fellowships (No. 101149710). The authors acknowledge the support of French-Swiss SMYLE Network. We thank Julien Joly (UFR ST, Dep Elec, UMLP) for 3D printed parts.